OpenStudy (jojomojo):

How can I get the answer 24 by only using the numbers 8,8,3,3. You can use the main signs add, subtract multiply and divide.

OpenStudy (anonymous):

8*3 just dont use the other 2

OpenStudy (homeworksucks):

8*3+(3-3)

OpenStudy (jojomojo):

1.8/8=1 2.1+3=4 3.4*3! 4*6(since 3 factorial=6) 4.{[(8/8)+3]*3!}=24.☺

OpenStudy (homeworksucks):

my way was simpler- no?

OpenStudy (anonymous):

\[\sqrt{8*8*3*3}\]

OpenStudy (jojomojo):

very much so

OpenStudy (tkhunny):

That's a tough one. There is only one distinct solution! 8/(3-(8/3)) @homeworksucks You way doesn't work. Extra 3 and one short on the 8s. @jojomojo Factorial is illegal. @ChmE Square Root is illegal.

OpenStudy (homeworksucks):

Oh right, derp again

OpenStudy (jojomojo):

aw man D:

OpenStudy (jojomojo):

Stranger wins this time